OpenStudy (annipuppi):

2|x-5|+3=5 show work

OpenStudy (annipuppi):

|2x+3|=3x+7

OpenStudy (annipuppi):

|5x-1|>=2

OpenStudy (tester97):

Can we start with one at a time?

OpenStudy (annipuppi):

Yeah start with the first one

OpenStudy (tester97):

OK so we have \[2|x-5|+3=5\] First step is subtracting 2 on both sides \[2|x-5|=2\] Now we can divide by 2 on both sides so we are just left with this \[|x-5|=1\] Since this is absolute value there is a special way we have to set up the equation to solve for x. We have to include both the positive and the negative so it will look something like this \[x-5=1 ~~or~-x+5=-1\] Do you think you can solve from here?

OpenStudy (annipuppi):

x=6 or x=4

OpenStudy (tester97):

Yes

OpenStudy (annipuppi):

Can I get help with the rest of them?